Suppose C and D represent two different school populations where C > D and C and D must be greater than 0. Which of the following expressions is the largest? Explain why. Show all work necessary.

A: (C + D)2

B: 2(C + D)

C: C2 + D2

D: C2 − D2

(10 points)


1
Expert's answer
2021-11-09T16:02:03-0500
"C>D>0"

"(C+D)^2=C^2+2CD+D^2>C^2+D^2"



"(C+D)^2=(C+D)(C+D)>(C+D)(C-D)"

"=C^2-D^2"

"(C+D)^2-2(C+D)=(C+D)(C+D-2)"

Therefore:

If "C+D>2," then


"A: (C + D)^2"

If "C+D<2," then


"B: 2(C + D)"